OpenStudy (bbb911):

i always give medals! ____x + 8____ + ____10 - 4x____= x + 7 x + 7 Add and simplify

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

is this a fraction

OpenStudy (anonymous):

lmao

OpenStudy (anonymous):

idek

OpenStudy (bbb911):

this is algebra 2 their factors

OpenStudy (anonymous):

can you write the problem using the equation editot?

OpenStudy (anonymous):

but where did the = sign come from

OpenStudy (anonymous):

\[\frac{ x+8 }{ x+7 }+\frac{ 10-4x }{ x+7 }\]

OpenStudy (anonymous):

ahh ok so thats it

OpenStudy (anonymous):

WIth equivalent denominators you can add the numerators and place them under that similar denominator.

OpenStudy (anonymous):

\[\frac{ 3(6 - x) }{ x + 7 }\]

OpenStudy (anonymous):

Giving the answer is not 'helping'. Be aware of that.

OpenStudy (anonymous):

i was about to ask her if she wanted the steps

OpenStudy (anonymous):

When we have same denominators then we just combine numerator & write same denominator at onces that is x+ 8 +10 - 4x / x+ 7 just simplify it

OpenStudy (anonymous):

add and then factor

OpenStudy (anonymous):

the silence is awkward.

OpenStudy (bbb911):

yes i would like the steps sorry i took long i was eating :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

l\\[\frac{ x+8 }{ x+7 }+\frac{ 10-4x }{ x+7 }\]

OpenStudy (bbb911):

ok step one is that then?

OpenStudy (anonymous):

\[factor 10 - 4x:\frac{ 10-4x }{ x+8 }=\frac{ 2(5-2x) }{ x+8 }\]

OpenStudy (bbb911):

so eliminate x + 7

OpenStudy (anonymous):

i messed on the denominator is x+7

OpenStudy (anonymous):

the denominator remains, i just focused on focused on factoring 10 - 4x

OpenStudy (bbb911):

oh ok so it factored out to 2(5 -2x) right

OpenStudy (bbb911):

from the 10 - 4x

OpenStudy (anonymous):

add \[\frac{ 2(5-5x) }{ x+7 } + \frac{ x+8 }{ ?\x+7 }\]

OpenStudy (anonymous):

yeah 10-4x is factored to 2(5-2x)

OpenStudy (anonymous):

my internet is mad slow

OpenStudy (bbb911):

its okay, im patient :) so when adding do i add across?

OpenStudy (anonymous):

2(5-2x) + (x + 8) = 3(6-x)

OpenStudy (anonymous):

yeah you go ahead and add across

OpenStudy (bbb911):

ok im a bit confused when adding , i kno you add across but to get 3 (6-x) im trying to figure tht out

OpenStudy (anonymous):

remember when I said factor 10 - 4x?

OpenStudy (bbb911):

yes

OpenStudy (anonymous):

omg still on this

OpenStudy (anonymous):

yeah scratch that, all i had to do was: (10 - 4x)+(x + 8) add 10 plus 8 18 - 4x + x add -4x + x 18 - 3x then factor

OpenStudy (anonymous):

my internet is slow its my fault

OpenStudy (anonymous):

factor to 3(6 - x)

OpenStudy (bbb911):

its ok :)

OpenStudy (bbb911):

oh i see thank you :)!!

OpenStudy (anonymous):

np